### v1

@@ 6 Conclusion
Hyper-differential sensitivity analysis with respect to model discrepancy has considerable potential to enable approximate optimization with quantified uncertainties for many applications where currently only low-fidelity optimization is possible. However, to provide useful uncertainty estimates, it is crucial to specify the model discrepancy prior distribution appropriately. For instance, if the model discrepancy prior distribution is not sufficiently informative, then the posterior will have large uncertainty in directions not informed by the high-fidelity data. Consequently, propagating this posterior discrepancy through the optimization problem will yield large uncertainty in the optimization solution such that it fails to be useful.
@@ Acknowledgements.
This work was supported by the Laboratory Directed Research and Development program at Sandia National Laboratories, a multimission laboratory managed and operated by National Technology and Engineering Solutions of Sandia LLC, a wholly owned subsidiary of Honeywell International Inc. for the U.S. Department of Energy’s National Nuclear Security Administration under contract DE-NA0003525. This written work is authored by employees of NTESS. The employees, not NTESS, own the right, title, and interest in and to the written work and is responsible for its contents. Any subjective views or opinions that might be expressed in the written work do not necessarily represent the views of the U.S. Government. The publisher acknowledges that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this written work or allow others to do so, for U.S. Government purposes. The DOE will provide public access to results of federally sponsored research in accordance with the DOE Public Access Plan.
@@ Appendix B Proof of Theorem 2
Furthermore, ${\tilde{{\bm{z}}}}^{T}{\bm{M}}_{\bm{z}}{\bm{W}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}}{\tilde{{\bm{z}}}}=||{\bm{E}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}}{\tilde{{\bm{z}}}}||_{{\bm{M}}_{\bm{z}}}^{2}$, and hence
$\lim\limits_{{\mathcal{U}}_{h}\to{\mathcal{U}},{\mathcal{Z}}_{h}\to{\mathcal{Z}}}{\tilde{{\bm{z}}}}^{T}{\bm{M}}_{\bm{z}}{\bm{W}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}}{\tilde{{\bm{z}}}}=\lim\limits_{{\mathcal{Z}}_{h}\to{\mathcal{Z}}}||{\bm{E}}_{\bm{z}}^{-1}{\bm{M}}_{\bm{z}}{\tilde{{\bm{z}}}}||_{{\bm{M}}_{\bm{z}}}^{2}=\|\mathcal{E}_{z}^{-1}\tilde{z}\|_{{\mathcal{Z}}}^{2},$
@@ Appendix E Empirical correlation length estimation
15:  Note: For some $i$ in lines 10-12, $x_{i}\pm\Delta\kappa\notin\Omega$, and hence evaluating $\hat{f}$ is not appropriate. These datapoints are omitted and the normalization factor $m-1$ in Line 12 is adjusted accordingly.

### v2

@@ 6 Conclusion
Hyper-differential sensitivity analysis with respect to model discrepancy has considerable potential to enable approximate optimization with quantified uncertainties for many applications where currently only low-fidelity optimization is possible. However, to provide useful uncertainty estimates, it is crucial to specify the model discrepancy prior distribution appropriately. As was demonstrated in Subsection 5.1, misspecification of the model discrepancy prior may lead to an optimal solution posterior distribution that is too narrow, thus giving unwarranted confidence in the solution. On the other hand, if the model discrepancy prior distribution is not sufficiently informative, then the posterior will have large uncertainty in directions not informed by the high-fidelity data. Consequently, propagating this posterior discrepancy through the optimization problem will yield large uncertainty in the optimization solution such that it fails to be useful.
@@
Traditional Bayesian approaches to hyper-parameter selection may be considered as alternatives to our approach. For instance, choosing hyper-parameters that maximize the data likelihood. However, there are several important problem characteristics that motivate our approach. We focus on problems where the number of high-fidelity model evaluations $N$ is small. In the case where $N=1$, there is no information in the data to inform $\alpha_{\bm{z}}$ using data likelihood maximization. For $N>1$ but much less than the dimension of $\bm{z}$}, there will be many local extrema in the data likelihood maximization as a result of the hyper-parameters having confounding effects. On the other hand, for any $N\geq 1$, our initialization approach leverages the physical interpretation of the hyper-parameters to sequentially select the hyper-parameters. The only solution variability comes from estimating expected values. Hybrid approaches can also be explored. For example, our proposed method can be combined with data likelihood maximization to (i) reduce the number of local minima by fixing a subset of hyper-parameters, (ii) inform the prior distribution of the hyper-parameters, or (iii) serve as an initialization for the data likelihood optimization algorithm.
@@ Acknowledgements.
This work was supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research Field Work Proposal Number 23-02526 (Ab-initio Visualization for Innovative Science). Li, Ouermi, and Johnson were partially supported by the Intel OneAPI CoE. Sandia National Laboratories is a multi-mission laboratory managed and operated by National Technology & Engineering Solutions of Sandia, LLC (NTESS), a wholly owned subsidiary of Honeywell International Inc., for the U.S. Department of Energy’s National Nuclear Security Administration (DOE/NNSA) under contract DE-NA0003525. This written work is authored by an employee of NTESS. The employee, not NTESS, owns the right, title and interest in and to the written work and is responsible for its contents. Any subjective views or opinions that might be expressed in the written work do not necessarily represent the views of the U.S. Government. The publisher acknowledges that the U.S. Government retains a non-exclusive, paid-up, irrevocable, world-wide license to publish or reproduce the published form of this written work or allow others to do so, for U.S. Government purposes. The DOE will provide public access to results of federally sponsored research in accordance with the DOE Public Access Plan. All data used in this article was generated by sampling from Gaussian random fields and solving partial differential equations. The source code is not publicly available, but data may be shared upon reasonable request to the corresponding author.
@@ Appendix B Proof of Theorem 2
[…]
@@ Appendix E Empirical correlation length estimation
15: Note: For some $i$ in lines 10-12, $x_{i}\pm\Delta\kappa\notin\Omega$, and hence evaluating $\hat{f}$ is not appropriate. These datapoints are omitted and the normalization factor m1m-1 in Line 12 is adjusted accordingly.